You want to save for retirement. Assuming you are now 20 years old and you want to retire at age 60, you have 40 years to watch your investment grow. You decide to invest in the stock market, which has earned about 8% per year over the past 80 years and is expected to continue at this rate. You decide to invest $2,000 today.
Answer #1

Future Value = Present Value * ( 1 + ( rate/100)) ^ ( time)

FV = 2000 * ( 1 + ( 8/100)) ^ 40 = $43,449.04

Excel Formula =FV(8%,40,0,-2000,0)

= $43,449.04